A body projected vertically up crosses a point `P` in its path at the end of 2 seconds and the highest point at the end of 5 seconds. After how many seconds from the start will it reach `P` again?

Grade:11

2 Answers

Rituraj Tiwari
3 years ago
Total Time taken for the ball to go upwards is (12+4)sec= 16 seconds...

And total time taken for the ball to go upwards and come down is(16+16)sec= 32 sec

As/question, It is asked for finding the time for coming back to p point so we need tosubtract (the whole time taken - q points time)=(32-12)second= 20 seconds..

The ball needs20 secondto reach the p point again
Harsh Chaudhari
one year ago
The body reaches point P at end of 2 second ...Therefore at point P t=3 starts
At maximum height , 5th second ends...Therefore at max height t=6 starts.
Time taken by body to reach max height from ground = 6s
Time taken to reach max height froh P = 3s
Time taken by body to go from P to max height is equal to time taken by body to descend from max height to point P.
Hence , 6+3 = 9s 
So the body reaches P again at the start of 9th seconds that is after 8 seconds are completed

Hence final answer is 8s.